Find the flux through the top,left,and right side of a cube of side 0.5 m when placed in a horizontal uniform electric field of 8.0N/C directed to the right.


1
Expert's answer
2022-02-13T17:52:17-0500

Electric flux


"\phi=E.A=EAcos\theta"

Where "\theta=90°"

"E=8N\/C"

"A=0.025m^2"

Top side

"\phi_t=E.A=EAcos\theta\\\theta=90\\\phi_t=0"

Bottom side

"\phi_b=E.A=EAcos\theta\\\theta=90\\\phi_b=0"

Right side

"\phi_r=EAcos\theta\\\phi_r=EAcos90=0"

Electric flux top left right side


"\phi_{net}=\phi_t+\phi_b+\phi_r+\phi_L+\phi_f+\phi_b"

"\phi=8\times0.25\times cos90°"

"\phi=0"